OpenStudy (anonymous):

Am I right? A cylinder labeled X has a height of 10 inches and a radius of 15 inches. Another cylinder labeled Y has a height of 40 inches and a radius of 60 inches. The volume of cylinder Y is _______ times the volume of cylinder X. The answer is 1.06

OpenStudy (nincompoop):

WHY MIGHT YOU THINK THAT YOU'RE WRONG?

OpenStudy (anonymous):

O-O....

OpenStudy (anonymous):

I think I'm wrong, because I'm not sure if I used the proper formula...?

OpenStudy (nincompoop):

THEN FIGURE OUT THE CORRECT FORMULA. THE INTERNET IS AVAILABLE FOR YOUT TO USE

OpenStudy (aravindg):

@nincompoop caps off please

OpenStudy (anonymous):

I tried searching already but failed to find it;;

OpenStudy (nincompoop):

ARAVIND, THERE'S NO RULE AGAINST USING UPPER CAPS.

OpenStudy (anonymous):

@Nurali , can you help with this problem?

OpenStudy (aravindg):

Though it seems like you are shouting

OpenStudy (aravindg):

@TheHero given any figure volume is got by the product of base area and height .Remember!

OpenStudy (aravindg):

Base of a cylinder is a circle :Arepi *r*r

OpenStudy (aravindg):

*Area=pi*r*r

OpenStudy (nincompoop):

http://bit.ly/13QMj74

OpenStudy (aravindg):

height=h

OpenStudy (aravindg):

So volume=pi *r*r*h

OpenStudy (anonymous):

3.14*15*15= 707.14 Volumi: 3.14*15*15*10= 7071.428

OpenStudy (anonymous):

Just divide the volume of Y by the volume of X. Pi x 60 x 60 x 40 / Pi x 15 x 15 x 10. If the final answer is 1.06 then ur correct. Y : X = 1.06 : 1 Which, by law of proportions means Y = 1.06 X; try it out

OpenStudy (anonymous):

3.14*60*60*40=452571.428

OpenStudy (anonymous):

;~; I did get 1.06 previously, I wasn't sure if it was right though.

OpenStudy (aravindg):

why not ?

OpenStudy (anonymous):

@AravindG somehow i think 1.06 is not right..radius itself is 4 times or something;

OpenStudy (anonymous):

D: Because, I just guessed the formula!

OpenStudy (aravindg):

If you get it by proper calculation .Why doubt it ?

OpenStudy (aravindg):

Oh ok ....Now I gave you the correct formula .Calculate again?

OpenStudy (anonymous):

I got 64.

OpenStudy (anonymous):

Yes i think Y is 64 times of X's volume..i think its correcto

OpenStudy (aravindg):

you are right !

OpenStudy (aravindg):

\[\huge \color {red}{\checkmark}\]

OpenStudy (anonymous):

Thanks so much! :DD

OpenStudy (aravindg):

yw :)

OpenStudy (anonymous):

yeah welcome